# The Southern Photometrical Local Universe Survey (S-PLUS): searching for metal-poor dwarf galaxies

### v1

@@ 4.1 Emission line fluxes and intensities
The observed fluxes of the detected emission lines, ), normalized to (H) and multiplied by 100, the extinction coefficients , and the extinction-corrected intensities normalized to are given in Table 2. The corresponding values for target ID 6, derived from the SDSS spectrum, are provided in the Appendix (Table 5). The uncertainties in the measured fluxes of each line are calculated using the expression from Tresse et al. (1999)
where is the mean standard deviation per pixel of the continuum determined on each side of the emission line, is the B600 spectral dispersion (0.1 nm for 2x binned spectra and 0.05 nm for unbinned spectra), is the number of pixels covered by the line, which we equated to the FWHM of the fitted Gaussian profile, and is the equivalent width of the line.
@@ 4.2 BPT diagram
4.3 Oxygen and nitrogen abundances from the direct method
@@
We determined the physical conditions of the star forming regions with the direct method (Pagel et al., 1992; Pilyugin & Thuan, 2005; Osterbrock & Ferland, 2006). Calculations of both the density and temperature of the electron gas were carried out using the Nebular Empirical Analysis Tool (NEAT, Wesson et al., 2012).
1 | 2 | 3 | 4 | 5a | 6 | |
Ne [cm-3] | 10857 | |||||
Te [K] | 18168908 | 166001900 | ||||
O/H | 2.240.02 | 3.29 | ||||
N/H | 0.750.04 | 1.93 | ||||
12+log(O/H) | 7.41 | 7.93 | 7.64 | 7.77 | 7.350.04 | 7.51 |
log(N/O) | -0.93 | -1.39 | -1.25 | -1.25 | -1.470.05 | -1.23 |
a Guseva et al. (2009).
@@
The electron temperature sensitive emission line [Oiii]4363 is detected in all the galaxies (Fig. 4), including the two additional objects with previous spectroscopic observations (ID 5 and 6). Temperatures and densities were derived using an iterative process from the relevant diagnostic lines with NEAT. We used the ratio of the [S ii]67176731 lines to estimate the electron densities, . We found that the ratio is around or above 1.4 for all galaxies, implying that all star-forming regions are in the low-density regime (Osterbrock & Ferland, 2006). Therefore we adopt a constant electron density value of 100 cm-3. The electron temperature ([O iii]) was calculated from the [O iii]4363/(4959 + 5007) emission-line intensity ratio and it was used to derive the abundances of O++ and O+. ranges between 14 000 K and 20000 K. Total oxygen abundances , were obtained from the intensities of the [O ii]73207330 and [O iii] lines. Although not relevant for O/H, additional ionization states that are present in the gas, but not in the optical region of the spectrum, are accounted for with ionization correction factors (ICFs). The ICF used to determine the nitrogen abundance (Kingsburgh & Barlow, 1994) is based on the only bright line detected in the observed wavelength range, [N ii]6584. Uncertainties are propagated from line flux measurements to the derived abundances with a Monte Carlo method implemented by NEAT.
@@
The oxygen abundances for target No. 5 were taken from Guseva et al. (2009), while for galaxy No. 6, we applied the same method as described above, using the [O iii] and [O ii] lines that were observed in the GMOS spectra (see Table 5). We find oxygen abundances varying between 12 + log(O/H) (between 5% and 17% of the solar value), confirming that the galaxies are metal poor. Electron temperatures, electron densities, oxygen and nitrogen abundances are displayed in Table 3.
@@ 4.4 The nitrogen-to-oxygen ratio
The nitrogen-to-oxygen ratio, N/O, is considered an indicator of the chemical evolutionary age of a galaxy. While oxygen is mostly formed by primary nucleosynthesis in massive stars ( M⊙) and injected into the ISM by core-collapse supernovae (Nomoto et al., 2013), nitrogen production is more complex and it can have both a primary and secondary origin. Primary nitrogen synthesis can occur in stars with masses around or above 1.3 M⊙ as part of the CNO cycle of oxygen and carbon during hydrogen burning. However stellar evolution models predict that it is mostly produced in intermediate-mass stars (4 - 8 M⊙) and released in the ISM during the asymptotic giant branch phase (AGB, Renzini & Voli, 1981; Woosley & Weaver, 1995; Meynet & Maeder, 2002). At low metallicities, nitrogen is produced from carbon and oxygen created via the helium burning process; in this case, nitrogen behaves like a primary element, its abundance being proportional to that of the other primary heavy elements resulting in a fixed ratio of N/O, independently of the oxygen abundance. Izotov & Thuan (1999) identified that the average value of the N/O ratio of extremely metal-poor galaxies (12 + log(O/H) 7.6) is approximately constant with a very small dispersion (N/O) , the so-called N/O plateau (Fig. 6). This was interpreted as the evidence of nitrogen being produced as a primary element by massive stars at low O/H. If the seed oxygen and carbon of the CNO cycle are produced in previous-generation stars, nitrogen synthesis is called secondary. At high metallicities (12 + log(O/H) ), the ISM metal abundance becomes high enough that the production of secondary nitrogen and N/O increases linearly with oxygen abundance (Tinsley, 1980; Pagel, 1985).
@@
More recent studies found that galaxies with low oxygen abundances show an excess N/O (van Zee & Haynes, 2006; Skillman et al., 2013; Watanabe et al., 2024), indicating that other factors can affect the ratio at low O/H. For example, strong winds carrying nitrogen-enriched gas from post-main sequence hot and massive Wolf-Rayet stars can raise N/O (Walsh & Roy, 1989; Maeder & Meynet, 2001; López-Sánchez et al., 2007; Telles et al., 2014). The increase in the dispersion could also be explained if the galaxy is going through a post-starburst phase: after the production of carbon and primary nitrogen by massive stars during the last major event of star formation, there is a delayed additional contribution of primary nitrogen produced by intermediate-mass stars in the post-starburst phase, unbalanced by the oxygen release of core-collapse supernovae (Garnett, 1990; van Zee & Haynes, 2006; Annibali et al., 2019). Another possibility could be provided by the rapid infall of metal-poor gas from the intergalactic medium that mixes with the ambient metal-rich ISM (Köppen & Hensler, 2005), diluting the oxygen abundance (i.e., it increases H in O/H). As a result, the mixed gas will be characterized by an unusually large N/O ratio for a given O/H (Sánchez Almeida et al., 2016b; Luo et al., 2021). An additional source of dispersion in the N/O versus O/H relation is the fact that oxygen can also be produced – through the third dredge-up in AGB stars – in low-metallicity galaxies, mainly for 12+log(O/H) 7.7, as was proved to happen in the ISM of the irregular dwarf galaxies of the Local Group (Magrini & Gonçalves, 2009).
@@
In Fig. 6, we show the N/O ratios of our targets compared to samples of star-forming galaxies from the literature with different ranges of metallicity (Izotov & Thuan, 1999; Izotov et al., 2006; van Zee & Haynes, 2006; Annibali et al., 2019). This figure shows the N/O plateau identified by Izotov & Thuan (1999) in extremely metal-poor galaxies (orange crosses), although more recent studies, also displayed in the figure, present a larger scatter at low metallicities raising the question about the actual existence of a plateau in this regime (van Zee & Haynes, 2006).
@@
We measured the nitrogen abundances and derived the N/O ratios for the four galaxies observed with GMOS and for target N. 6, and we used the values derived in Guseva et al. (2009) for the remaining dwarf (Table 3). We find (N/O) : three out of six dwarfs have typical values given their metallicity range and the scatter observed in other samples, while the remaining three show a higher than expected N/O. Particularly, ID 1 is the galaxy with the largest N/O (log(N/O) = -0.93), which would be more typical of a galaxy with a ten times larger metallicity. We did not detect the He ii 4686 line, which would suggest the presence of Wolf-Rayet stars in any of the dwarfs.
@@ 4.5 H line profile fitting and star formation rates
To derive the star formation rates of our targets, we extracted 1D spectra using larger apertures than those employed to estimate the oxygen abundances, in order to include the entire H emission detected in the GMOS spectra (see Fig. 3). The aperture sizes vary between 78 and 137 (Tab. 4). Figure 7 shows the region of the spectra centered around the H line, obtained for these larger apertures. H fluxes, displayed in Table 4, were obtained modeling the line profiles with PySpecKit.
For most of the targets, a satisfactory fit is achieved using two Gaussian components, except for ID 4, where three Gaussians are needed (Fig. 7). As can be seen in the fourth bottom panel of Fig. 3, this galaxy hosts a brighter component with a wider velocity dispersion (knot 3), and two dimmer ones exhibiting narrower (knots 1 and 2). The kinematical centers of the lower luminosity narrow components are offset from those of the main knot, producing the net effect of an overall asymmetry in the global line profiles. Knot 1 is on the blue side of the spectra (approaching gas), while knot 2 is on the red side of the spectra (receding gas). Despite being separated by only approximately 120 pc, knot 1 appears blueshifted at high velocities compared to knot 2. The offset in wavelength between the two star-forming clumps is , corresponding to a velocity difference of km s-1(bottom-right panel in Fig. 7). This might suggest that knot 1 is kinematically distinct from the rest of the galaxy, and further investigation is needed to better address this issue more thoroughly.
@@
Overall, metal-poor galaxies are found to have starburst-level SFRs; however, our sample shows normal star formation activities with SFRs ranging from 0.01 to 0.1 M⊙ yr-1 and specific star formation rates (sSFRs) around 10-10 yr-1. Figure 8 compares sSFRs and stellar masses of our sample with those estimated by H18 and L23. These two samples cover a wider range of these parameters. Particularly, the selection criteria of H23 are such that their sample is more biased towards higher redshift, more active metal-poor galaxies. The solid line shows the star formation sequence defined by Schiminovich et al. (2007) in the SDSS as a reference for the overall trend between the two parameters.
@@
4.6 Mass-metallicity relation
@@
The presence of metal-poor outliers at a given stellar mass is in general attributed to external events. This can happen if star formation occurs in pockets of pristine gas accreted from the intergalactic medium (Sánchez Almeida et al., 2016a; De Rossi et al., 2017; Ju et al., 2022; del Valle-Espinosa et al., 2023), or if it is triggered by a merger/interaction of a close galaxy pairs (Ellison et al., 2008b; Lelli et al., 2012; Grossi et al., 2020). In these scenarios, the metallicity of the ISM will be diluted and the galaxy would appear as an outlier in the MZR relation as well as in the luminosity-metallicity relation (LZR). Such an effect would be even stronger in the LZR (Ekta & Chengalur, 2010; McQuinn et al., 2020) due to the high luminosity of young massive stars and the contribution of strong emission lines, making the star-forming dwarfs brighter in the band when compared to galaxies of similar mass with a lower star formation activity.
Outflows due to supernova feedback ejecting enriched gas from the star-forming regions could generate metal losses in low-mass systems (McQuinn et al., 2015, 2019). However, given the low star formation rates observed, it is less likely that strong outflows due to stellar feedback may be causing the observed metal deficiency (Grossi et al., 2020).
@@ 5 Summary and conclusions
1. All the objects with spectroscopic follow-up were found to be metal-poor galaxies, with 12 + log(O/H) , i.e., between 5% and 17% of the solar abundance, confirming the effectiveness of our method based on color cuts and multi-band SED fitting.
2. We estimated nitrogen abundances and nitrogen-to-oxygen ratios. Half of the galaxies have N/O ratios compatible with the observed metallicity, while three dwarfs show excess N/O for a given O/H. Particularly, ID 1, the most metal-deficient target of the sample, exhibits a N/O more typical of a galaxy with ten times higher metallicity.

### v2

@@ 4.3 Extinction correction
The observed fluxes of the detected emission lines, ), normalized to (H) and multiplied by 100, the extinction coefficients , and the extinction-corrected intensities normalized to are given in Table 2. The corresponding values for target ID 6, derived from the SDSS spectrum, are provided in the Appendix (Table 5).
@@ 4.4 BPT diagram
4.5 Oxygen and nitrogen abundances
@@
We determined the physical conditions of the star forming regions with the direct method (Pagel et al., 1992; Pilyugin & Thuan, 2005; Osterbrock & Ferland, 2006). Calculations of both the density and temperature of the electron gas were carried out using NEAT (Wesson et al., 2012).
| 1 | 2 | 3 | 4 | 5 | 6 | |
| Ne [cm-3] | 920200 | |||||
| Te [K] | 16600 700 | 18400900 | 156001000 | |||
| O+/H | 5.83 | 4.88 | 6.15 | 8.42 | 9.38 | 8.51 |
| O++/H | 3.05 | 6.13 | 3.33 | 4.34 | 0.94 | 2.79 |
| O/H | 1.88 | 3.65 | ||||
| ICF(O) | 1 | 1 | 1 | 1 | 1 | 1 |
| N+/H | 6.16 | 4.82 | 4.68 | 7.93 | 3.71 | 6.40 |
| N/H | 0.740.07 | 2.75 | ||||
| ICF(N) | 6.2 0.7 | 13.7 1.6 | 6.4 0.5 | 2.000.05 | 4.30.5 | |
| 12+log(O/H) | 7.56 0.06 | 7.82 | 7.59 | 7.71 0.05 | 7.280.05 | 7.56 0.08 |
| log(N/O) | -0.98 0.07 | -1.00 | -1.12 | -1.02 | -1.400.07 | -1.12 |
@@
The auroral line [Oiii]4363 is detected in all the galaxies (Fig. 5), including the two additional objects with previous spectroscopic observations (ID 5 and 6). Temperatures and densities were derived using an iterative process from the relevant diagnostic lines with NEAT. We used the ratio of the [S ii]67176731 lines to estimate the electron densities, . For all GMOS targets the ratio is close to or slightly above 1.4 for all galaxies, indicating that the star-forming regions lie in the low-density regime (Osterbrock & Ferland, 2006). We therefore ran NEAT assuming a constant electron density value of 100 cm-3. This assumption has a negligible effect on the derived oxygen abundances, since the direct-method abundances are primarily sensitive to (Osterbrock & Ferland, 2006). The electron temperature ([O iii]) was calculated from the [O iii]4363/(4959 + 5007) emission-line intensity ratio and it was used to derive the abundances of O++ and O+. ranges between 14 000 K and 20000 K. Total oxygen abundances , were obtained from the intensities of the [O ii]73207330 and [O iii] lines. The contribution of oxygen ionization states not emitting in the optical region is negligible for the total oxygen abundance (Izotov et al., 2006; Andrews & Martini, 2013). NEAT computes the oxygen ionization correction factor, ICF(O), following Kingsburgh & Barlow (1994): ICF(O) = (O/H)/[(He+/He++]/He+)2/3. In practice, this factor is very close to unity, since both O+ and O++ are observed in our spectra.
For nitrogen, Table 2 and Fig. 5 show that only [N II]6584 is detected. NEAT therefore estimates the ionic abundance (N+/H+) from this line, and converts it into the total nitrogen abundance using ICF(N)=(O/H)/(O+/H+) (Kingsburgh & Barlow, 1994). For both N+ and O+ ionic abundances, we adopt Te[O III]. The more appropriate low-ionization temperature (derived from N+, O+, or S+ lines) cannot be used, since the auroral lines of these species are not detected (see Table 2). Lastly, the N/O abundance ratio is derived directly from N+/O+, in order to avoid relying on the ICF required to obtain the total nitrogen abundance. As noted above, ICF(N) is necessarily larger than ICF(O), that is essentially equal to unity. Uncertainties are propagated from line flux measurements to the derived abundances with a Monte Carlo method implemented by NEAT. Table 3 presents the total abundances, together with the ionic abundances and the ICFs used to derive the total values.
@@
For target No. 5, Guseva et al. (2009) reported an oxygen abundance of 12 + log(O/H) = 7.35 0.04. For consistency with our analysis, we used the extinction corrected emission line intensities from their work and recalculated the oxygen and nitrogen abundances with NEAT finding 12 + log(O/H) = 7.28 0.05 (Table 5). Galaxy No. 6 was observed by the SDSS; we extracted the emission line fluxes as described above and derived the oxygen abundance using the same [O iii] and [O ii] lines that were observed in the GMOS spectra (see Table 5). We find oxygen abundances in the range 12 + log(O/H) (i.e. between 4% and 13% of the solar value), confirming that the galaxies are metal poor. Electron temperatures, electron densities, oxygen and nitrogen abundances, ICF factors are displayed in Table 3.
@@ 4.6 The nitrogen-to-oxygen ratio
The nitrogen-to-oxygen ratio, N/O, is considered an indicator of the chemical evolutionary age of a galaxy. While oxygen is mostly formed by primary nucleosynthesis in massive stars ( M⊙) and injected into the ISM by core-collapse supernovae (Nomoto et al., 2013), nitrogen production is more complex and it can have both a primary and secondary origin. Secondary nitrogen synthesis can occur in stars with masses around or above 1.3 M⊙ as part of the CNO cycle of oxygen and carbon during hydrogen burning. However stellar evolution models predict that it is mostly produced in intermediate-mass stars (4 - 8 M⊙) and released in the ISM during the asymptotic giant branch phase (AGB, Renzini & Voli, 1981; Woosley & Weaver, 1995; Meynet & Maeder, 2002). At low metallicities, nitrogen is produced from carbon and oxygen created via the helium burning process; in this case, nitrogen behaves like a primary element, its abundance being proportional to that of the other primary heavy elements resulting in a fixed ratio of N/O, independently of the oxygen abundance. Izotov & Thuan (1999) identified that the average value of the N/O ratio of extremely metal-poor galaxies (12 + log(O/H) 7.6) is approximately constant with a very small dispersion (N/O) , the so-called N/O plateau (Fig. 7). This was interpreted as the evidence of nitrogen being produced as a primary element by massive stars at low O/H. At high metallicities (12 + log(O/H) ), the ISM metal abundance becomes high enough that the production of secondary nitrogen and N/O increases linearly with oxygen abundance (Tinsley, 1980; Pagel, 1985; Vila-Costas & Edmunds, 1993; Pilyugin et al., 2012; Andrews & Martini, 2013; Vincenzo et al., 2016).
@@
More recent studies found that galaxies with low oxygen abundances show an excess N/O (van Zee & Haynes, 2006; Skillman et al., 2013; Watanabe et al., 2024; Zinchenko et al., 2024), indicating that other factors can affect the ratio at low O/H. For example, strong winds carrying nitrogen-enriched gas from post-main sequence hot and massive Wolf-Rayet stars can raise N/O (Walsh & Roy, 1989; Maeder & Meynet, 2001; López-Sánchez et al., 2007; Telles et al., 2014). The increase in the dispersion could also be explained if the galaxy is going through a post-starburst phase: after the production of carbon and primary nitrogen by massive stars during the last major event of star formation, there is a delayed additional contribution of primary nitrogen produced by intermediate-mass stars in the post-starburst phase, unbalanced by the oxygen release of core-collapse supernovae (Garnett, 1990; van Zee & Haynes, 2006; Annibali et al., 2019). Another possibility is the rapid infall of metal-poor gas from the intergalactic medium that mixes with the ambient more metal-rich ISM (Köppen & Hensler, 2005). This process dilutes the oxygen abundance (by increasing the hydrogen content in O/H), without significanlty altering the N/O ratio. As a result, the mixed gas will be characterized by an apparent excess in the N/O – O/H plane (Sánchez Almeida et al., 2016b; Luo et al., 2021).
Although probably small, another source of dispersion in the N/O versus O/H relation is the fact that oxygen can also be produced in low-metallicity galaxies through the third dredge-up in AGB stars, mainly at 12+log(O/H) 7.7. This has been observed in the ISM of Local Group irregular dwarf galaxies (Magrini & Gonçalves, 2009). However, it is not possible to quantify this effect in galaxies where only H ii regions are detected, as in our study.
@@
In Fig. 7, we show the N/O ratios of our targets compared to samples of star-forming galaxies from the literature with different ranges of metallicity (Izotov & Thuan, 1999; Izotov et al., 2006; van Zee & Haynes, 2006; Andrews & Martini, 2013; Annibali et al., 2019; Zinchenko et al., 2024) and the model of Vila-Costas & Edmunds (1993). This figure shows the N/O plateau identified by Izotov & Thuan (1999) in extremely metal-poor galaxies (plus symbols), although more recent studies, also displayed in the figure, present a larger scatter at low metallicities raising the question about the actual existence of a plateau in this regime (van Zee & Haynes, 2006; Zinchenko et al., 2024).
@@
We measured the nitrogen abundances and derived the N/O ratios for the six galaxies of our sample. We find (N/O) : five out of six dwarfs show a higher than expected N/O. Particularly, ID 1 is the galaxy with the largest N/O (log(N/O) = -0.98), which would be more typical of a galaxy with a ten times larger metallicity. We did not detect the He ii 4686 line, which would suggest the presence of Wolf-Rayet stars in any of the dwarfs. However, a detailed analysis of the N/O ratio and its dependence on oxygen abundance is beyond the scope of this paper. Further observations are required to detect additional nitrogen lines, which would allow a more reliable determination of the oxygen abundances in these galaxies.
@@ 4.7 H line profile fitting and star formation rates
To derive the star formation rates of our targets, we extracted 1D spectra using larger apertures than those employed to estimate the oxygen abundances, in order to include the entire H emission detected in the GMOS spectra (see Fig. 4). The aperture sizes vary between 78 and 137 (Tab. 4). Figure 8 shows the region of the spectra centered around the H line, obtained for these larger apertures. H fluxes, displayed in Table 4, were obtained by integrating the line profiles after subtracting the continuum with PySpecKit (Ginsburg et al., 2022).
@@
Although metal-poor galaxies can have starburst-level SFRs (Izotov et al., 2019, 2024b), our sample shows normal star formation activities with SFRs ranging from 0.01 to 0.1 M⊙ yr-1 and specific star formation rates (sSFRs) around 10-10 yr-1. Figure 9 compares sSFRs and stellar masses of our sample with those estimated by H18 and L23. Stellar masses have been derived using the same methodology for all three samples, i.e., from the r - i colour described in Eq. 1. While the H18 stellar masses were already estimated in this way, we have recalculated the stellar masses of the L23 sample using the SDSS photometry provided in this work. These two samples cover a wider range of these parameters. Particularly, the selection criteria of H23 are such that their sample is more biased towards higher redshift, more active metal-poor galaxies. The solid line shows the star formation sequence defined by Schiminovich et al. (2007) in the SDSS as a reference for the overall trend between the two parameters.
@@
4.8 Mass-metallicity and luminosity-metallicity relations
@@
The presence of metal-poor outliers at a given stellar mass is in general attributed to external events. This can happen if star formation occurs in pockets of pristine gas accreted from the intergalactic medium (Sánchez Almeida et al., 2016a; De Rossi et al., 2017; Ju et al., 2022; del Valle-Espinosa et al., 2023), or if it is triggered by a merger/interaction of a close galaxy pairs (Ellison et al., 2008b; Lelli et al., 2012; Grossi et al., 2020). In these scenarios, the metallicity of the ISM will be diluted and the galaxy would appear as an outlier in the MZR relation as well as in the luminosity-metallicity relation (LZR). Such an effect would be even stronger in the LZR (Ekta & Chengalur, 2010; McQuinn et al., 2020) due to the high luminosity of young massive stars and the contribution of strong emission lines, making the star-forming dwarfs brighter in the band when compared to galaxies of similar mass with a lower star formation activity. In the right panel of Fig. 10 we display the luminosity metallicity relation of all the samples. Following L23 we derive the luminosity of the galaxies in the -band using the empirical relation of Cook et al. (2014) based on and magnitude:
The galaxies are compared to the LVL survey and, as expected, the offset from the main relation is even more prominent in the luminosity-metallicity relation (right panel of Fig. 10). Outflows due to supernova feedback ejecting enriched gas from the star-forming regions could generate metal losses in low-mass systems (McQuinn et al., 2015, 2019). However, given the low star formation rates observed, it is less likely that strong outflows due to stellar feedback may be causing the observed metal deficiency (Grossi et al., 2020).
@@ 5 Summary and conclusions
1. All the objects with spectroscopic follow-up were found to be metal-poor galaxies, with 12 + log(O/H) , i.e., between 4% and 13% of the solar abundance, confirming the effectiveness of our method based on color cuts and multi-band SED fitting.
2. We estimated nitrogen abundances and nitrogen-to-oxygen ratios. Five out of six galaxies show excess N/O for a given O/H. Particularly, ID 1, the most metal-deficient target of the sample, exhibits a N/O more typical of a galaxy with ten times higher metallicity.